Quantitative Problem: Bellinger Industries is considering two projects for inclusion in its capital budget, and you have been asked to do the analysis Both projects after-tax cash flows are shown on the time line below. Deprediation, salvage values, net operating working capital requirements, and tax effects are all included in these cash flows. Both projects have 4-year lives, and they have risk characteristics similar to the firms average project. Bellingers WACC is 8%. 2 3 4 290 440 340 Project A-1,500 Project B1,500 700 300 370 305 790 What is Project As MIRR? Do not round intermediate calculations. Round your answer to two decimal places. olo Show All Feedback What is Project Bs MIRR? Do not round intermediate calculations. Round you answer to two decimal places.
Answer #1

A:

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.
Hence future value of inflows=700*(1.08)^3+370*(1.08)^2+290*(1.08)+340

=$1966.5664

MIRR=[future value of inflows/Present value of outflows]^(1/time period)-1

=[$1966.5664/$1500]^(1/4)-1

=7.01%(Approx).

B:

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.
Hence future value of inflows=300*(1.08)^3+305*(1.08)^2+440*(1.08)+790

=$1998.8656

MIRR=[future value of inflows/Present value of outflows]^(1/time period)-1

=[1998.8656/$1500]^(1/4)-1

=7.44%(Approx).